Compute the mean and standard deviation of these numbers: 1, 5, 3, 1. Hint: The sum of squares for the sample is 11. Use two decimal places. Mean: and Standard Deviation:
Answer #1

Given that ,

Data : 1, 5, 3, 1

The sum of squares for the sample is 11

n = 4

Mean = = = (1 + 5+ 3 + 1) / 4 = 10 / 4 = 2.5

standard deviation = s = (x - )2 / n - 1  

= 11 / 3

= 1.91485